if a mole were to contain 1*1024 particle what would be the mass of:: 
(a)one mole of oxygen .
(b)a single oxygen molecule.

Grade:12th pass

3 Answers

Arun
5 years ago

1) Mass of 6.022 x 1023 particles of oxygen = 32 g/mole

  Mass of 1024 particles of oxygen = (32/6.022 x 1023 ) x 1024

   = 53.13 g/mole

 

Thus, mass of one mole of oxygen = 53.13 g

 

2) Mass of one molecule of oxygen = 53.13/1024

   = 53.13 x 10-24 g

Rihaan Vashistha
3 years ago
Here consider 1 mole having 1×10^24 particles1) mass of one mole of oxygen= molecular mass
Jack
3 years ago
Mass of 6.023 ×1023 perticale of oxygen =32 gm/mol .mass of 1×1024 perticale  =(1×1024×32)÷(6.023×1023)=53.13gm